# Nanoscale Quantum Imaging of Spin Dynamics using a Hybrid 2D/3D System

### v2

@@ Discussion
The crystal flakes were grown by the atmospheric pressure high temperature (APHT) method, as previously described in detail [47, 48]. Briefly, the process starts by mixing high purity elemental boron, enriched in the 10B isotope to 99%, with nickel and chromium with weight ratios of 4:48:48, respectively. The mixture is then heated under a blanket of nitrogen enriched in the 15N isotope to 99.5% at a pressure of 850 torr to 1550°C, to produce a homogeneous molten solution. After 24 hours, the solution is slowly cooled at 4°C/h, to 1500°C, then at 50°C/h to 1300°C, and 200°C/h to room temperature. The solubility is decreased as the temperature is reduced, causing crystals to precipitate on the surface of the metal. The flakes were exfoliated from the metal with thermal release tape. The free-standing crystalline flakes were typically 10 to 20 m thick. To create boron vacancies, the flakes were neutron-irradiated. Samples were irradiated in the Ohio State University nuclear reactor, operated at a power 300 kW for 1 hour, corresponding to neutron fluences of neutron/cm2.
Photoluminescence (PL) and Raman spectroscopy were performed using a commercial InVia Qontor confocal system (Renishaw) to characterize defect implantation. The sample was excited using a 532 nm laser delivered through a Leica objective (NA = 0.85). Scattered signals were collected by the same objective, spectrally filtered using ultra-narrow notch filters, and subsequently dispersed onto diffraction gratings with groove densities of 1800 lines mm-1 (Raman measurements) and 300 lines mm-1 (PL measurements).
@@
Continuous-wave (CW) optically detected magnetic resonance (ODMR) measurements of the NV center and the ensemble, as shown in Figure 2b and 4, were performed under continuous laser excitation and microwave drive. A static magnetic field was applied using a neodymium ring magnet mounted on a three-axis translation stage, while the microwave frequency was swept. To suppress fluorescence from the ensemble, a 750 nm short-pass filter was used during NV ODMR measurements.
Pulsed NV relaxometry measurements were conducted using a 3 s laser pulse to initialize the NV spin into the state. The spin was then allowed to relax toward thermal equilibrium during a dark interval , during which microwave excitation was applied. Relaxation was detected as a reduction in photoluminescence (PL) during a subsequent readout sequence, preceding the next initialization pulse. At each readout pulse sequence, a reference PL signal, PL(0), is recorded after re-polarizing the spin into the state. By varying , an exponential decay curve was obtained, from which the value was extracted.
@@
Acknowledgments
Funding:
@@ Acknowledgments
H.Z. conceived the project and performed the NV relaxometry experiments. A.L.M. and H.Z. analyzed the data, and co-wrote the manuscript with input from all other authors. A.L.M., Y.W., R.G. and G.H. developed the theoretical model, with additional input from L.L., C.Z., and J.T.D. R.G., G.H. and C.Z. performed the Monte Carlo simulation and quantified the defect density. I.V.V. synthesized the hBN samples and performed optical spectroscopy. T.P. and J.H.E. synthesized the samples and organized the neutron irradiation. S.R. conducted the helium ion implantation. S.G. fabricated the coplanar waveguide. Y.C.W. and B.J.L. assisted with the microwave delivery setup. A-P.L., S.J., and B.J.L. provided technical support for the scanning NV microscope measurements.

### v3

@@ Discussion
The crystal flakes were grown by the atmospheric pressure high temperature (APHT) method, as previously described in detail [32, 26]. Briefly, the process starts by mixing high purity elemental boron, enriched in the 10B isotope to 99%, with nickel and chromium with weight ratios of 4:48:48, respectively. The mixture is then heated under a blanket of nitrogen enriched in the 15N isotope to 99.5% at a pressure of 850 torr to 1550°C, to produce a homogeneous molten solution. After 24 hours, the solution is slowly cooled at 4°C/h, to 1500°C, then at 50°C/h to 1300°C, and 200°C/h to room temperature. The solubility is decreased as the temperature is reduced, causing crystals to precipitate on the surface of the metal. The flakes were exfoliated from the metal with thermal release tape. The free-standing crystalline flakes were typically 10 to 20 m thick. To create boron vacancies, the flakes were neutron-irradiated. Samples were irradiated in The Ohio State University nuclear reactor, operated at a power 300 kW for 1 hour, corresponding to neutron fluences of neutron/cm2. Prior to the cross-relaxometry experiments, micrometer-scale flakes were exfoliated from irradiated macroscopic crystals using standard mechanical exfoliation. This process yielded flakes with thicknesses of approximately 50 nm, used for the measurements shown in Fig. 2, and approximately 15 nm, used for the data presented in Fig. S26.
Photoluminescence (PL) and Raman spectroscopy were performed on hBNnat using a commercial InVia Qontor confocal system (Renishaw) to characterize defect implantation. The sample was excited using a 532 nm laser delivered through a Leica objective (NA = 0.85). Scattered signals were collected by the same objective, spectrally filtered using ultra-narrow notch filters, and subsequently dispersed onto diffraction gratings with groove densities of 1800 lines mm-1 (Raman measurements) and 300 lines mm-1 (PL measurements). The PL spectrum of the samples was not measured but is expected to have a similar spectrum reported in previous studies [19].
@@
Continuous-wave (CW) optically detected magnetic resonance (ODMR) measurements of the NV center and the ensemble, as shown in Figure 1c and 3, were performed under continuous laser excitation and microwave drive. A static magnetic field was applied using a neodymium ring magnet mounted on a three-axis translation stage, while the microwave frequency was swept. To suppress fluorescence from the ensemble, a 750 nm short-pass filter was used during NV ODMR measurements.
Pulsed NV relaxometry measurements were conducted using a 3 s laser pulse to initialize the NV spin into the state. The spin was then allowed to relax toward thermal equilibrium during a dark interval , during which microwave excitation was applied. Relaxation was detected as a reduction in photoluminescence (PL) during a subsequent readout sequence, preceding the next initialization pulse. At each readout pulse sequence, a reference PL signal, PL(0), is recorded after re-polarizing the spin into the state. By varying , exponential decay curves such as those shown in Fig. 1d were obtained and fit, from which the value was extracted.
@@
Data availability
All data needed to evaluate the conclusions of this paper are available in the Article or the Supplementary Information. The experimental and simulation data generated in this study are provided in the Source Data file. All raw data relevant to this work are available from the corresponding author upon reasonable request. Source data are provided with this paper.
Acknowledgments:
@@
H.Z. conceived the project and performed the NV relaxometry experiments. A.L.M. and H.Z. analyzed the data, and co-wrote the manuscript with input from all other authors. A.L.M., Y.W., R.G. and G.H. developed the theoretical model, with additional input from L.L., C.Z., and J.T.D. R.G., G.H. and C.Z. performed the Monte Carlo simulation and quantified the defect density. I.V.V. synthesized the hBN samples and performed optical spectroscopy. T.P. and J.H.E. synthesized the samples and organized the neutron irradiation. S.R. conducted the helium ion implantation. S.G. fabricated the coplanar waveguide. Y.C.W. and B.J.L. assisted with the microwave delivery setup. A.-P.L., S.J., and B.J.L. provided technical support for the scanning NV microscope measurements.
@@
Materials & Correspondence:
Correspondence and requests for materials should be addressed to Huan Zhao